OpenStudy (anonymous):

sin25° = a and cos40°=b Express the following in terms of a and b (1-tan155° sin295° ) / (tan410° cos 230° )

OpenStudy (jamesj):

You've got it. Great.

OpenStudy (anonymous):

yoo

OpenStudy (anonymous):

I just click wrongly :(

OpenStudy (anonymous):

\(\tan 155 = \tan(180-25) =\cdots \)

OpenStudy (jamesj):

Yes, for each of the trig terms, you want to figure out how to write them as differences like this from standard trig functions. For example, with the one above: because tan(180) = 0 and tan(-x) = -tan(x), it follows that tan(180-25) = [ tan(180) - tan(25) ] / [ 1 - tan(180).tan(-25) ] = - tan(25) = - sin(25)/sqrt(1-sin^2(25)) = - a/sqrt(1-a^2)